# Can LLM-Augmented autonomous agents cooperate?, An evaluation of their cooperative capabilities through Melting Pot

### v1

@@ 4.2 Cognitive modules
Insight Generation: The second stage involves utilizing these questions to retrieve pertinent memories from long-term memory. Subsequently, the questions and retrieved memories are employed to generate three insights, which are then stored as reflections in the long-term memory.
The retrieval of relevant memories employs a weighted average encompassing cosine similarity, recency score, and poignancy scores. The recency score is computed as , where denotes the number of hours since the last memory was recorded. Meanwhile, the poignancy score reflects the intensity assigned to the memory at its point of creation. Throughout the experiments, a uniform poignancy score of 10 was assigned to all memory types. For the complete prompt and more details on this module, refer to Appendix E.
Action Module: This module plays the role of generating an action for the agent to undertake. As detailed in Appendix F, the selection of the action is determined by the Language Model (LLM), which considers the agent’s comprehension of the world, its current goals and plans, reflections, ongoing observations, and the available valid actions within the environment. The creation of new action sequences occurs under two conditions: when the current sequence is empty or when the agent is responding to observations. For this prompt, we manually crafted a reasoning structure, similar to those described in Self-Discover [12] to help the LLM consider different alternatives, and evaluate them before making the final decision.
@@ 5 Evaluation scenarios
To assess the outcomes, we utilized the per capita average reward of the focal population as our primary metric. The focal population comprises LLM agents, and the chosen metric aligns with the Melting Pot framework’s approach [2], which evaluates population welfare. We compare this metric across two sets of scenarios.
The first set of scenarios has the intention of measuring how the personality given to the agents affects the agents’ welfare. For this purpose, we prepared five scenarios: (1) as a baseline we do not give the agents any personality specifications (Without personality), (2) we tell the agents to be cooperative (All coop.), (3) we tell the agents to be cooperative and provide a short description of how to be cooperative in the chosen scenario (All coop. with def.), (4) we tell the agents to be selfish (All selfish), (5) we tell the agents to be selfish and provide a definition with the expected behavior of someone selfish for the given scenario (All selfish with def.).
The second set of scenarios is more challenging as the competition increases by reducing the number of trees and apples, modifying the agents’ initial social environment understanding, or by adding other entities to the environment (bots), these changes demand a deeper understanding from the agents and swift reactions to master the scenarios. More concretely, the • first three scenarios consist of an environment where there are three agents and only one apple tree. Each scenario differs in the personality given to the agents: (1) All coop, (2) All selfish, and (3) without personality. The last scenario of the second set (4) has the same base configuration, but there are two agents and two bots, where the bots are reinforcement learning agents that were trained to harvest in an unsustainable way and to attack other agents. These bots make part of scenario 0 of the commons harvest open scenario described in Meltingpot 2.0 [2].
We also add a scenario aimed at demonstrating how the information an agent has about the rest of the agents can influence their behavior. In this scenario, the environment begins with the same number of trees; however, from the start of the simulation, each agent was informed that among them, one was acting entirely selfishly, representing a risk due to their unsustainable consumption.
@@
Ten simulations for each scenario were done in which the LLM agents were powered by the GPT3.5 from the OpenAI API for the majority of modules, and GPT4 powered the act module. On another hand, the Ada model was used to create contextual embeddings of the memories. A detail of the simulations cost is available in the Appendix G.
@@ 6.1 Impact of personality in population welfare
The average per capita reward obtained for the first set of scenarios is shown in Fig. 3. The best-performing simulations were the ones where no particular personality description was given to the agents, followed by the scenarios where the agents were instructed to be selfish. Surprisingly, the scenarios where the agents were told to be cooperative had the worst performance. On further analysis, we found that these results are explained mainly by the number of times the agents decided to attack other agents (see Fig. 4).
To gain a better understanding of the agents’ behavior, we recorded the number of times the agents decided to attack other agents, and the number of times these attacks were effective. These actions play an important part in the game because they are the only mechanism the game provides to directly interact with other agents. Therefore, helping the agents counteract other agents’ behavior, such as attacking other agents that are taking apples indiscriminately to avoid the depletion of the apple trees or decreasing the competition when there are too many agents near the same tree. More concretely, when an agent attacks and the ray beam hits its target (another agent), the agent that was hit is taken out of the game for the next five steps and then it is revived in a random position of the spawning area of the map.
In Fig. 4 we show the results of these attack indicators for the first set of experiments. The results depict some important differences across the scenarios, mainly reflecting the reluctance of the cooperative agents to attack, and a strange difference between the number of attacks of the selfish agents with definition and the selfish agents without definition.
@@
Another important behavior to keep track of is the decisions the agents made when they were near the last apple of a tree. Whether they choose to take it or ignore it is a crucial event and highly impactful in the final per capita reward, as there are only 6 apple trees in the game, and taking the last apple of a tree would mean that the tree would be depleted and will not produce more apples. For this reason, we created an indicator that counts how many times the agents closed the distance between the last apple of a tree and its position, divided by how many times the nearest apple to the agent was the last apple of a tree. However, this indicator does not consider that sometimes the last apple, despite being the closest to the agent, is not visible to the agent because it is outside the observation window of the agent. This limitation could have had an impact on the observed results.
In Fig. 5 we can see that the proportion of times the agents moved towards the last apple is pretty similar across all the scenarios, indicating that the personality descriptions did not cause a major effect on the awareness of the agents regarding the welfare detriment caused by the depletion of apple trees. These results highlight a limited understanding among the agents regarding the consequences of their actions.
@@ 6.2 Performance of the agents in more challenging scenarios
The second set of experiments consists of scenarios where the competence increases or the resources become scarcer. The purpose of these scenarios is to measure how the agents respond to the new game conditions.
@@
One single tree scenarios. The first three scenarios in this set represent an environment involving a more intensive competition for resources. The three agents, who usually have a limited field of vision, are in constant observation of a single tree in the environment, which is situated in a confined space. The difference between each scenario lies in the type of personality assigned to each agent, with the personalities in this case being All cooperative, All selfish, and Without Personality. For practical purposes, no specific definition was given to any personality. The purpose of the scenario is to demonstrate the collective sustainability capacity that different types of agents can have where resources are highly limited.
@@
In Fig. 6, the results for the ”Per capita reward” contrasted with the ”Average amount of available apples” for the described group of scenarios are observed. Upon close examination, it is noted that the slope of the reward curve for cooperative agents is less than that of the ”Selfish” and ”Without personality” agents, this behavior contributes to this set of agents having resource availability for a slightly longer period, as shown in the figure. However, given the dynamics of the probability of apple reappearance, this behavior was not significant enough to allow cooperative agents to have a considerably superior reward per capita. Therefore, it is concluded that any set of agents was able to demonstrate sufficiently good sustainable behavior due to their lack of understanding of the world and their lack of communication and coordination capabilities with other agents.
@@
Agents versus Bots. The fifth scenario of the second set of experiments exposes two agents to the presence of two reinforcement learning bots. The policy of the bots makes them take the apples without regard for the replenishment rate or the risk of depleting the trees, they focus solely on maximizing their rewards by taking the apples, but they also attack other agents, mainly where there are no other apples in proximity.
@@
In Fig. 7 we get the results of the average reward per capita for the agents versus bots scenario. The initial notable observation is that the bots can consistently achieve higher rewards than the agents. This phenomenon is mainly explained by the policy of the bots that prioritize taking all the visible apples over other actions, while the agents explore the map or go to other positions of the map with higher frequency than the bots. However, it is important to note how the per capita reward for the bots stops increasing earlier than that for the agents, indicating greater difficulty for the bots to increase their rewards when trees are scarce in comparison to the agents. Moreover, we found that in half of the simulations, at least one of the agents achieved a better reward than that of a bot, leading us to conclude that sometimes the agents are capable of performing better than the greedy policy of the agents. By closer examination, we observed that in those simulations, the agents were able to find apple trees more easily than the bots and that they also tended to attack when another agent or bot was taking apples from the same tree as them.
@@
Moreover, Fig. 9 shows that the agents depleted trees with higher frequency than the agents. Thus the agents demonstrated the capacity to sometimes restrain themselves from just taking apples by trying to maximize their long-term rewards, whereas bots always prioritized their short-term reward.
@@ 6.3 Impact of knowledge of other agent’s behavior
This experiment considers the hypothetical scenario in which all agents are previously informed that one specific agent is entirely selfish and the implications that its uncooperative behavior can have. Likewise, this agent is informed to act selfishly (providing the previously described definition of selfishness). The objective of this scenario is to highlight the behavior that agents can exhibit when possessing valuable information about their social environment.
@@ 7.1 Importance of Cooperative Capabilities
To address the proposed scenarios in a better way, agents needed to recognize certain principles. For instance, they should refrain from harvesting the last apple in a green patch to prevent depletion and should engage in cooperation with other agents while avoiding collaboration with the bots or uncooperative agents. Observing that the bots consistently harvested apples unsustainably, agents should have deduced that attacking the bots was necessary to protect the green patches from depletion. This ability to prioritize long-term and collective welfare over short-term rewards, as well as recognizing the divergent behavior and preferences of other entities (bots), aligns with what Dadfoe et al. [1] refer to as cooperative capabilities.
This prompts a consideration of whether current agent architectures genuinely enable cooperative behavior, and if the absence of such capabilities hinders their ability to navigate more intricate tasks and environments. Dadfoe et al. [1] succinctly categorize cooperative capabilities into four essential components:
@@
In essence, cultivating collaborative capabilities within agent architectures is crucial for tackling the complexities inherent in diverse tasks and environments. Historically, agent architectures have inadequately endowed agents with such capabilities. Instances such as Generative Agents [9] and the Improving Factuality and Reasoning in Language Models through Multiagent Debate [3] enable agents to engage in conversations or observe the perspectives of others. However, these approaches are hampered by the absence of independent evaluation criteria and discernment specific to the current LLMs limitations.
@@ 7.2 Cooperative Agent Architecture
Based on previous findings, we proposed an architecture to enhance agents’ cooperative capabilities (see Fig. 11). In this architecture several new modules are proposed:
@@
Understanding module: This component is tasked with a comprehensive analysis of the agent’s memories, fostering a deeper comprehension of the surrounding world. The agent’s proficiency extends to predicting the behaviors of fellow agents and discerning environmental changes, enabling it to take actions with a keen awareness of their potential consequences. Notably, the agent must possess the capacity to infer both the governing principles of the world and the underlying motivations guiding others’ actions. This inference capability extends to scenarios where these principles may deviate from common knowledge or the pre-training model knowledge. Zhu et al. [13] demonstrate that LLMs, like GPT4, can learn such rules when explicitly prompted to identify them, utilizing question-answer pairs to later apply the learned rules in problem-solving. The proposed module operates by initially extracting the rules and behavioral patterns of the world and other agents. It achieves this by prompting the LLM with historical world observations and the current state of the world, aiming to identify rules that explain the current state based on the agent’s observations. These identified rules are initially stored as world hypotheses. As the agent utilizes these hypotheses to interpret the current state, they are transformed into explicit rules once they surpass a predefined threshold. Additionally, the LLM is prompted to generate predictions about future states of the environment, empowering the agent to make informed decisions guided by anticipated future scenarios.
@@ 8 Conclusion and Future Work
Cooperative capabilities have been somewhat overlooked in LLMs agent architectures, yet they may represent the crucial element enabling agents to accomplish pioneering tasks and thrive in intricate environments. As large language models (LLMs) advance, agent architectures stand to gain significantly by attaining enhanced responses from LLMs, particularly in tasks demanding substantial reasoning or when confronted with copious information in the prompt.
In this paper, our objective is to ascertain whether LLMs-enhanced autonomous agents can operate cooperatively. To this end, we adapt the Melting Pot scenarios to textual representations that can be easily operationalized by LLMs, and implement a reusable architecture for the development of LAAs employing the modules proposed in [9]. This architecture includes short and long-term memories and cognitive modules of perception, planning, reflection, and action. The common “Commons Harvest” game was used to test the resulting system, and the results were evaluated from the viewpoint of cooperative metrics in different proposed scenarios.
The results indicate a gap in the current agent’s cooperative capabilities vis-à-vis unfamiliar situations. Agents showed a cooperative tendency but lacked adequate understanding of how to collaborate effectively in an unknown environment. The agents needed to understand complex factors like the need to conserve resources, identify non-cooperative agents, and prioritize collective welfare over short-term gains. The research, thereby, draws attention to the need for a more inclusive architecture fostering cooperation and enhancing agent capabilities, including superior understanding, effective communication, credible commitment, and well-defined social structures or institutions.

### v2

@@ 4.2.3 Reflection module
Insight Generation: The second stage involves using these questions to retrieve pertinent memories from long-term memory. Subsequently, the questions and retrieved memories are employed to generate three insights, which are then stored as reflections in the long-term memory.
The retrieval of relevant memories employs a weighted average encompassing cosine similarity, recency score, and poignancy scores. The recency score is computed as , where denotes the number of hours since the last memory was recorded. Meanwhile, the poignancy score reflects the intensity assigned to the memory at its point of creation. Throughout the experiments, a uniform poignancy score of 10 was assigned to all memory types. For the complete prompts and more details on question formulation and insight generation processes of this module, refer to Appendix E.
4.2.4 Action Module
This module plays the role of generating an action for the agent to undertake. As detailed in Appendix F, the selection of the action is determined by the Language Model (LLM), which considers the agent’s comprehension of the world, its current goals and plans, reflections, ongoing observations, and the available valid actions within the environment. The creation of new action sequences occurs under two conditions: when the current sequence is empty or when the agent is responding to observations. For this prompt, we manually crafted a reasoning structure, similar to those described in Self-Discover (Zhou et al., 2024), to help the LLM consider different alternatives and evaluate them before making the final decision.
@@ 5 Evaluation scenarios
To assess the outcomes, we utilized the per capita average reward of the focal population as our primary metric. The focal population comprises LLM agents, and the chosen metric aligns with the Melting Pot framework’s approach (Agapiou et al., 2023), which evaluates population welfare. We compare this metric across two sets of scenarios.
The first set of scenarios is intended to measure how the personality given to the agents affects their welfare. For this purpose, we prepared five scenarios: (1) as a baseline we do not give the agents any personality specifications (Without personality), (2) agents are instructed to be cooperative (All coop.), (3) agents are instructed to be cooperative and provide a short description of how to be cooperative in the chosen scenario (All coop. with def.), (4) agents are instructed to be selfish (All selfish), (5) agents are instructed to be selfish and provide a definition with the expected behavior of someone selfish for the given scenario (All selfish with def.).
The second set of scenarios is more challenging as competition increases by reducing the number of trees and apples, modifying the agents’ initial understanding of the social environment, or by adding other entities to the environment (bots). These changes demand a deeper understanding from the agents and swift reactions to master the scenarios. More concretely, the • first three scenarios consist of an environment where there are three agents and only one apple tree. Each scenario differs in the personality given to the agents: (1) all cooperative, (2) all selfish, and (3) without personality. The last scenario of the second set (4) has the same base configuration, but with two agents and two bots, where the bots are reinforcement learning agents trained to harvest unsustainably and attack other agents. These bots are part of scenario 0 of the commons harvest open scenario described in Meltingpot 2.0 (Agapiou et al., 2023).
We also add a scenario aimed at demonstrating how the information an agent has about the rest of the agents can influence their behavior. In this scenario, the environment starts with the same number of trees; however, from the beginning of the simulation, each agent is informed that among them, one is acting entirely selfishly, representing a risk due to their unsustainable consumption.
@@
Ten simulations for each scenario were conducted where the LLM agents were powered by the GPT-3.5 from the OpenAI API for the majority of modules, and GPT-4 powered the action module. On the other hand, the Ada model was used to create contextual embeddings of the memories. Details of the simulation costs are available in Appendix G.
@@ 6.1 Impact of personality in population welfare
The average per capita reward obtained for the first set of scenarios is shown in Fig. 3. The best-performing simulations were those where no particular personality description was given to the agents, followed by the scenarios where the agents were instructed to be selfish. Surprisingly, the scenarios where the agents were told to be cooperative had the worst performance. Further analysis revealed that these results are primarily explained mainly by the number of times the agents decided to attack other agents (see Fig. 4).
To gain a better understanding of the agents’ behavior, we recorded the number of times the agents decided to attack other agents, and the number of times these attacks were effective. These actions are crucial in the game as they are the only mechanism provided for direct interaction with other agents. They help agents counteract behaviors such as indiscriminate apple picking by other agents, which threatens the depletion of apple trees, or decreasing competition when too many agents are near the same tree. More concretely, when an agent attacks and the ray beam hits its target (another agent), the agent that was hit is taken out of the game for the next five steps and then revived in a random position of the spawning area of the map.
Fig. 4 shows the results of these attack indicators for the first set of experiments. The results depict some important differences across the scenarios, mainly reflecting the reluctance of the cooperative agents to attack, and an unexpected difference between the number of attacks of the selfish agents instructed with definition and the selfish agents without definition.
@@
Another important behavior to track is the decisions the agents made when they were near the last apple of a tree. Whether they choose to take it or ignore it is a crucial event and highly impactful on the final per capita reward, as there are only six apple trees in the game, and taking the last apple from a tree means that the tree would be depleted and would not produce more apples. For this reason, we created an indicator that counts how many times the agents closed the distance between themselves and the last apple of a tree, divided by how many times the nearest apple to the agent was the last apple of a tree. However, this indicator does not account for situations where the last apple, despite being the closest to the agent, is not visible to the agent because it is outside the observation window of the agent. This limitation could have impacted the observed results.
In Fig. 5, we can see that the proportion of times the agents moved towards the last apple is pretty similar across all the scenarios, indicating that the personality descriptions did not cause a major effect on the awareness of the agents regarding the welfare detriment caused by the depletion of apple trees. These results highlight a limited understanding among the agents regarding the consequences of their actions.
@@ 6.2 Performance of the agents in more challenging scenarios
The second set of experiments consists of scenarios where the competition increases or the resources become scarcer. The purpose of these scenarios is to measure how the agents respond to the new game conditions.
6.2.1 One single tree scenarios
@@
The first three scenarios in this set represent an environment involving a more intensive competition for resources. The three agents, who usually have a limited field of vision, are in constant observation of a single tree in the environment, which is situated in a confined space. The difference between each scenario lies in the type of personality assigned to each agent, with the personalities in this case being All cooperative, All selfish, and Without Personality. For practical purposes, no specific definition was given to any personality. The purpose of the scenario is to demonstrate the collective sustainability capacity that different types of agents can have where resources are highly limited.
@@
In Fig. 6, the results for the “Per capita reward” are contrasted with the “Average amount of available apples” for the described group of scenarios. Upon close examination, it is noted that the slope of the reward curve for cooperative agents is less than that for Selfish and Without personality agents. This behavior contributes to this set of agents having resource availability for a slightly longer period, as shown in the figure. However, given the dynamics of the probability of apple reappearance, this behavior was not significant enough to allow cooperative agents to have a considerably superior reward per capita. Therefore, it is concluded that no set of agents was able to demonstrate sufficiently good sustainable behavior due to their lack of understanding of the world and their lack of communication and coordination capabilities with other agents.
6.2.2 Agents versus Bots
@@
The fifth scenario of the second set of experiments exposes two agents to the presence of two reinforcement learning bots. The policy of the bots makes them take the apples without regard for the replenishment rate or the risk of depleting the trees; they focus solely on maximizing their rewards by taking the apples, but they also attack other agents, especially where there are no other apples in proximity.
@@
In Fig. 7 we see the results of the average reward per capita for the agents versus bots scenario. The initial notable observation is that the bots consistently achieve higher rewards than the agents. This phenomenon is mainly explained by the policy of the bots, which prioritizes taking all the visible apples over other actions, while the agents explore the map or move to other positions on the map with higher frequency than the bots. However, it is important to note how the per capita reward for the bots stops increasing earlier than that for the agents, indicating greater difficulty for the bots to increase their rewards when trees are scarce, compared to the agents. Moreover, we found that in half of the simulations, at least one of the agents achieved a better reward than that of a bot, leading us to conclude that sometimes the agents are capable of outperforming the greedy policy of the bots. Upon closer examination, we observed that in those simulations, the agents were able to find apple trees more easily than the bots and that they also tended to attack when another agent or bot was taking apples from the same tree as them.
@@
Moreover, Fig. 9 shows that the agents depleted trees with higher frequency than the bots. Thus, the agents demonstrated the capacity to sometimes restrain themselves from just taking apples by trying to maximize their long-term rewards, whereas bots always prioritized their short-term rewards.
@@ 6.3 Impact of knowledge of other agent’s behavior
This experiment considers the hypothetical scenario in which all agents are previously informed that one specific agent is entirely selfish and the implications that its uncooperative behavior can have. Likewise, this agent is informed to act selfishly, providing the previously described definition of selfishness. The objective of this scenario is to highlight the behavior that agents can exhibit when possessing valuable information about their social environment.
@@ 7.1 Importance of Cooperative Capabilities
To address the proposed scenarios in a better way, agents needed to recognize certain principles. For instance, they should refrain from harvesting the last apple in a green patch to prevent depletion and should engage in cooperation with other agents while avoiding collaboration with the bots or uncooperative agents. Observing that the bots consistently harvested apples unsustainably, agents should have deduced that attacking the bots was necessary to protect the green patches from depletion. This ability to prioritize long-term and collective welfare over short-term rewards, as well as recognizing the divergent behavior and preferences of other entities (bots), aligns with what Dafoe et al. (2020) refer to as cooperative capabilities.
This prompts a consideration of whether current agent architectures genuinely enable cooperative behavior, and if the absence of such capabilities hinders their ability to navigate more intricate tasks and environments. Dafoe et al. (2020) succinctly categorize cooperative capabilities into four essential components:
@@
In essence, cultivating collaborative capabilities within agent architectures is crucial for tackling the complexities inherent in diverse tasks and environments. Historically, agent architectures have inadequately endowed agents with such capabilities. Instances such as Generative Agents (Park et al., 2023) and the Improving Factuality and Reasoning in Language Models through Multiagent Debate (Du et al., 2023) enable agents to engage in conversations or observe the perspectives of others. However, these approaches are hampered by the absence of independent evaluation criteria and discernment specific to the current limitations of LLMs.
@@ 7.2 Cooperative Agent Architecture
Based on previous findings, we propose an architecture to enhance agents’ cooperative capabilities (see Fig. 11). In this architecture, several new modules are proposed:
@@
Understanding module: This component is tasked with a comprehensive analysis of the agent’s memories, fostering a deeper comprehension of the surrounding world. The agent’s proficiency extends to predicting the behaviors of fellow agents and discerning environmental changes, enabling it to take actions with a keen awareness of their potential consequences. Notably, the agent must possess the capacity to infer both the governing principles of the world and the underlying motivations guiding others’ actions. This inference capability extends to scenarios where these principles may deviate from common knowledge or the pre-training model knowledge. Zhu et al. (2023) demonstrate that LLMs, like GPT-4, can learn such rules when explicitly prompted to identify them, utilizing question-answer pairs to later apply the learned rules in problem-solving. The proposed module operates by initially extracting the rules and behavioral patterns of the world and other agents. It achieves this by prompting the LLM with historical world observations and the current state of the world, aiming to identify rules that explain the current state based on the agent’s observations. These identified rules are initially stored as world hypotheses. As the agent utilizes these hypotheses to interpret the current state, they are transformed into explicit rules once they surpass a predefined threshold. Additionally, the LLM is prompted to generate predictions about future states of the environment, empowering the agent to make informed decisions guided by anticipated future scenarios.
@@ 8 Conclusion and Future Work
Cooperative capabilities have been somewhat overlooked in LLMs’ agent architectures, yet they may represent the crucial element enabling agents to accomplish pioneering tasks and thrive in intricate environments. As large language models (LLMs) advance, agent architectures stand to gain significantly by attaining enhanced responses from LLMs, particularly in tasks demanding substantial reasoning or when confronted with copious information in the prompt.
In this paper, our objective is to ascertain whether LLMs-enhanced autonomous agents can operate cooperatively. To this end, we adapt the Melting Pot scenarios to textual representations that can be easily operationalized by LLMs, and implement a reusable architecture for the development of LAAs employing the modules proposed in Generative Agents (Park et al., 2023). This architecture includes short and long-term memories, and cognitive modules of perception, planning, reflection, and action. The “Commons Harvest” game was used to test the resulting system, and the results were evaluated from the viewpoint of cooperative metrics in different proposed scenarios.
The results indicate a gap in the current agents’ cooperative capabilities vis-à-vis unfamiliar situations. Agents showed a cooperative tendency but lacked an adequate understanding of how to collaborate effectively in an unknown environment. The agents needed to understand complex factors like the need to conserve resources, identify non-cooperative agents, and prioritize collective welfare over short-term gains. The research thereby draws attention to the need for a more inclusive architecture fostering cooperation and enhancing agent capabilities, including superior understanding, effective communication, credible commitment, and well-defined social structures or institutions.
@@
Data Availability
All the data generated for each simulation and the summary files for each experiment are available at experiments data. The code repository will be shared on GitHub upon acceptance of the paper.